OpenStudy (anonymous):

Help please? Aubrey uses her credit card to buy some clothes for $552.86. She can pay up to $195 on the credit card each month. What is the total cost of her purchase if the APR on the credit card is 27.3% compounded monthly? $552.86 $578.05 $703.79 $740.85

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

She makes the purchase for $552.86. After one month, she owes $552.86 + the interest of that month. One month's interest is 27.3%/12 on the balance, so $552.86 * 0.273/12 = $12.58 At the end of the first month, she owes $552.86 + $12.58 = $565.44. She pays $195. Now she owes $565.44 - $195 = $370.44 After the second month, she owes $370.44 + interst of that month. One month's interest is 27.3%/12 on the balance, so $370.44 * 0.273/12 = $8.43 At the end of the second month, she owes $370.44 + $8.43 = $378.87 She pays $195. Now she owes $378.87 - $195 = $183.87 After the third month, she owes $183.87 + interest of that month. One month's interest is 27.3%/12, so $183.87 * 0.273/12 = $4.18 At the end of the third month, she owes $183.87 + $4.18 = $188.05 She pays $188.05 and pays it off. The total amount she paid was $195 + $195 + $188.05 = $578.05 (this is the correct one lol)

OpenStudy (anonymous):

thank you :) how about this one? Taylor has a credit card with an interest rate of 24.5% compounded monthly that she must pay off in seven years. If her balance on the credit card is $2,345, what must her monthly payment be? $27.92 $58.61 $222.31 $732.52

OpenStudy (anonymous):

im answering in the other question